Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
4x – 3y = 1
-12x + 9y = 5

Answer by elima(1433)

4x – 3y = 1
-12x + 9y = 5
==================
(3)4x-3y=1
==================
12x+9y=3
-12x+9y=5
----------
No solution, both variables will be eliminated.
:)